# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

### v1

@@ 7.1 Constructing and utilizing multi-modal data
Considering report generation as a multi-modal problem is more aligned with clinical practice (Tu et al., 2024; Yan et al., 2023). Babar et al. (2021a) have proven the ineffectiveness of simple encoder-decoder report generation models and mentioned that adding prior knowledge can be a promising method. However, the current utilization of multi-modal data remains under-explored. Firstly, the methods for non-image feature extraction and the fusion of multi-modality data are often limited and simplistic, such as using graph encoding for the knowledge base and attention mechanisms for feature fusion. Secondly, the construction of the knowledge base is imperfect. The pre-defined graph (Zhang et al., 2020) is overly simplistic. Despite Radgraph being a vast knowledge base, it is solely derived from reports, lacking the relationship between images and reports, such as organ recognition or understanding of typical radiological scenarios, which radiologists possess. The Chest ImaGenome dataset (Wu et al., 2021) provides the organ recognition annotations, alleviating the problem. Additionally, as far as we know, publicly available knowledge bases only concentrate on chest X-rays, leaving a gap in general medical knowledge databases.
@@ 7.2 Evaluation of medical correctness
Evaluating the medical correctness of generated reports is crucial for the clinical application. Compared to previous works (Messina et al., 2022; Liao et al., 2023), recent works have paid more attention to it, but still have two shortcomings. First, in the reviewed articles, medical correctness evaluation has only been applied to chest X-ray reports. Secondly, the evaluation is based on the automatic labeler of radiology reports, which are only targeted at 14 types of diseases and the average F1 score is around 0.798 (Smit et al., 2020). Thus, improving the accuracy and scale of automatic labeling tools can help optimize the evaluation process.
7.3 Large public datasets and unified comparison benchmark
@@
As shown in Table 1, most of the public datasets are limited in size. Deep learning-based techniques require a large amount of data. The contemporary prevalence of large language models underscores this need for extensive data volumes. Among the datasets, the MIMIC dataset is relatively large but only includes Chest X-ray data. Large datasets targeting other image modalities and diseases need to be constructed. In addition, while MIMIC-CXR is a well-established benchmark compared to IU-XRay, dataset utilization lacks standardization, complicating comparisons. We urge papers using the MIMIC dataset to define their training, validation, and testing partitions, with explicit disclosure of the filtration method, particularly for the testing dataset.
7.4 Human-AI interaction
Most papers overlook the interaction between users (e.g., clinicians or patients) and automated systems. When the system performs as an AI assistant, users may want to know the insights of the model regarding specific aspects of medical images. In the reviewed works, Tanwani et al. (2022) constructed a Visual Question Answering system for medical report generation, and Tanida et al. (2023) linked the output results to image regions using object detection, allowing users to select areas of interest and receive corresponding language explanations. Recently, dialogue systems (e.g., GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), Gemini (Team et al., 2023)) based on large language models and large multi-modal models have shown people more possibilities for human-AI interaction. Although general large language models can provide answers to some questions in medical question answering benchmarks, their deployment in clinical settings remains unfeasible due to safety concerns within the medical domain (Yan et al., 2023; Saab et al., 2024).
To enable dialogue systems to comprehend medical knowledge, fine-tuning the model with medical data is an intuitive approach, such as LLaVA-Med (Li et al., 2024) and Med-PaLM 2 (Singhal et al., 2023). However, they did not test the model’s performance on the report generation task. Saab et al. (2024) proposed Med-Gemini, a series of highly proficient multi-modal models tailored specifically for the medical domain. They intuitively showed the interactive report generation process for a normal case, while lacking quantitative results to evaluate report generation performance. Exploring the direction of report generation with human-AI interaction holds significant promise. Additionally, fine-tuning large models demands substantial GPU resources, making efficient methods crucial.
@@ 7.5 Standardized report generation
Most of the existent methods focused on unstructured report generation, while structured reporting has several advantages, such as saving time (Hong and Kahn, 2013; Nobel et al., 2022), preventing errors, decreasing communication expenses linked to ambiguous natural language. Recently, Pellegrini et al. (2023) developed a structured template and released a related dataset. This promising start could pave the way for further exploration in this direction.
@@ 8 Conclusions
In this study, we have conducted a detailed technical review of 89 papers on automatic medical report generation published in the years 2021, 2022, 2023, and 2024 to showcase both mainstream and novel techniques. Our particular focus lies on the utilization and fusion of multi-modality data. The analysis of methods is structured based on the components of the report generation pipeline, presenting the key techniques for each component. Additionally, we provide a summary of the current publicly available datasets and evaluation methods, encompassing both quantitative and qualitative assessments. Subsequently, we compare the results from a subset of the 89 papers under the same experimental setting. Finally, we outline the current challenges and propose future directions for medical report generation. Overall, sustained progress is needed to produce standardized and clinically accurate reports. This survey aims to offer a comprehensive overview of report-generation techniques, emphasize critical issues, and assist researchers in promptly grasping recent advancements in the field to build more robust systems for clinical practice.
@@
Table A1: Summary of papers in the survey. The overview is based on the findings of the survey. The dataset and metrics section focuses on mainstream datasets and metrics. ’–’ indicates that the paper either does not detail this process or does not include the key techniques summarized in this survey. The following abbreviations are used: I-Archi: the architecture of image feature learning, I-Module: the enhancement module of image feature learning, NI: the feature learning of non-image data, GEI: gastrointestinal endoscope image, RetiI: retinal image, Term: terminology, KnowB: knowledge base, RealR: real report, ClinicalI: clinical information, Ques: questionnaires, FreFilter: frequency-based filtering, ConAtoL: converting all tokens to lowercase, RemoNAT: removing non-alphabetic tokens, AT: auxiliary task, ContrasL: Contrastive learning, MM: memory metric, FeatO: feature-level operation, OptimS: optimization strategies, H-LSTM: hierarchical LSTM, ReLoss: re-weighted loss function, ReinL: reinforcement learning, R-L: Rouge-L, C-D: CIDEr-D, CE: clinical efficacy, Com: comparison, Clas: classification, Escore: error scoring, ECI: extracting case-related information, IU: IU X-Ray, MIMIC: MIMIC-CXR, ImaGeno: Chest ImaGenome, COV: COV-CTR, DeepEye: DeepEyeNet, CCCT: Chinese COVID-19 CT, Ret-I: Retina ImBank, and Ret-C: Retina Chinese. In addition, AT-Graph, AT-Class, AT-EC, and AT-DS mean the graph-based, classification, embedding comparison, and detection/segmentation auxiliary tasks.
Paper Input data Data preparation Feature Learning Feature Fusion Generation Training Strategy Datasets Metrics
I-Archi I-Module NI
@@
ResNet AT-Graph Transformer FeatO Transformer IU,
@@
Liu et al. [2021d] Chest X-ray ResNet AT-Graph Transformer IU,
@@
CNNs LSTMs Curricu-
lum lea-
rning
IU,
@@
You et al. [2021] Chest X-ray ResNet Attention Transformer IU,
@@
Miura et al. [2021] Chest X-ray DenseNet+
Transformer
Transformer ReinL MIMIC BLEU,
@@
Alfarghaly et al. [2021] Chest X-ray Resizing DenseNet AT-Class Attention Transformer IU BLEU,
@@
BioSentVec
Attention H-LSTM IU,
@@
CNNs Embedding
layer
FeatO LSTM DeepEye BLEU,
@@
DenseNet AT-Class BERT FeatO,
Attention
Transformer AT CCCT BLEU,
@@
Endo et al. [2021] Chest X-ray ResNet Retrieval MIMIC BLEU,
@@
ResNet AT-Class LSTM H-LSTM ReinL IU,
@@
Pino et al. [2021] Chest X-ray DenseNet Template IU,
@@
ResNet ContrasL H-LSTM IU,
@@
DenseNet Transformer MIMIC BLEU,
@@
CNNs Embedding
layer
LSTM LSTM DeepEye BLEU,
@@
InceptionV3 AT-Class LSTM IU,
@@
FreFilter
ResNet RotatE,
ECI,
@@
DenseNet+
Transformer
AT-EC Transformer,
@@
DenseNet AT-Graph H-LSTM IU BLEU,
@@
Transformer
AT-Graph Transformer FFA-IR BLEU,
@@
Transformer
MM Transformer ReinL IU,
@@
Transformer
ContrasL MM Transformer IU,
@@
DenseNet+
Transformer
AT-Graph BERT FeatO,
Attention
Transformer IU BLEU,
@@
Transformer Transformer IU BLEU,
@@
Yan [2022] Chest X-ray DenseNet+
Transformer
AT-Graph,
MM
Transformer IU,
@@
Transformer
AT-Class Transformer AT IU BLEU,
@@
Transformer
AT-Class Attention,
@@
METEOR
Wang et al. [2022d] Chest X-ray Resizing Transformer AT-Class,
@@
CNNs Embedding
layer
Attention LSTM DeepEye BLEU,
@@
DenseNet ContrasL Attention Transformer IU,
@@
Jia et al. [2022] Chest X-ray DenseNet Attention H-LSTM IU,
@@
Zhang et al. [2022] Chest X-ray Resizing VGG Retrieval MIMIC Precision
@@
Transformer
AT-Graph Embedding
layer
Attention Transformer MIMIC BLEU,
@@
Transformer
AT Attention Transformer IU,
@@
DenseNet AT-Class LSTM ReLoss IU BLEU,
@@
METEOR
Abela et al. [2022] Chest X-ray DenseNet Template MIMIC BLEU,
@@
Wu et al. [2022] Chest X-ray ResNet ContrasL LSTM IU,
@@
ContrasL
BERT Attention Transformer IU BLEU
@@
METEOR
Kong et al. [2022] Chest X-ray Resizing Transformer Retrieval IU,
@@
Transformer AT-Class,
ContrasL
FeatO Transformer MIMIC BLEU,
@@
Wang et al. [2023d] Chest X-ray Tokenizing Transformer AT Transformer IU,
@@
DenseNet AT-Graph Attention Transformer IU,
@@
Transformer
Transformer ReinL,
@@
Tokenizing Transformer AT-Class,
ContrasL
ECI,
@@
FeatO,
Attention
Transformer IU,
@@
Zhang et al. [2023b] Chest X-ray DenseNet+
Transformer
@@
DenseNet+
Transformer
AT-Class ECI,
@@
Transformer
ContrasL Transformer IU,
@@
DenseNet+
Transformer
BERT FeatO,
Attention,
@@
DenseNet ContrasL Transformer FeatO,
Attention
Transformer IU,
@@
Pellegrini et al. [2023] Chest X-ray,
Ques
EfficientNet BERT Attention Classification - Rad-ReStruct CE
@@
AT-DS Embedding
layer
Attention Transformer MIMIC,
@@
Attention Transformer - IU,
@@
Embedding
layer
Attention Transformer - MIMIC,
@@
Jeong et al. [2024] Chest X-ray Resizing Transformer Retrieval MIMIC BLEU,
Escore
@@
Transformer FeatO,
Attention
Transformer ReinL IU,
@@
Transformer
AT Transformer ReLoss IU,
@@
Transformer
- Transformer,
Attention
@@
ResNet AT-Class Transformer FeatO,
Attention
Transformer IU,
@@
Table A2: Comparisons of the model performance on the MIMIC-CXR Dataset. B1, B2, B3, B4, R-L, C-D, P, R, and F represent BLEU-1, BLEU-2, BLEU-3, BLEU-4, ROUGE-L, CIDEr-D, precision, recall, and F1 score, respectively. The best and second best results are highlighted. All values were extracted from their papers.

### v2

@@ 9.1 Constructing and utilizing multi-modal data
Considering report generation as a multi-modal problem is more aligned with clinical practice (Tu et al., 2024; Yan et al., 2023). Babar et al. (2021a) have proven the ineffectiveness of simple encoder-decoder report generation models and mentioned that adding prior knowledge can be a promising method. However, the current utilization of multi-modal data remains under-explored. Firstly, the methods for non-image feature extraction and the fusion of multi-modality data are often limited and simplistic, such as using graph encoding for the knowledge base and attention mechanisms for feature fusion. Additionally, there is a lack of comparative analysis on the effectiveness of these methods. For instance, the reviewed papers do not compare two attention fusion approaches: one where concatenated features from different modalities are used as the input for Q, K, and V, and another where one modality is used for Q, and the other modality for K and V. Secondly, there is inadequate utilization of complementary information across input modalities and a lack of targeted design for filtering different types of noise. Thirdly, the construction of the knowledge base is imperfect. The pre-defined graph (Zhang et al., 2020) is overly simplistic. Despite Radgraph being a vast knowledge base, it is solely derived from reports, lacking the relationship between images and reports, such as organ recognition or understanding of typical radiological scenarios, which radiologists possess. The Chest ImaGenome dataset (Wu et al., 2021a) provides organ recognition annotations, alleviating the problem. Additionally, these publicly available knowledge bases only concentrate on chest X-rays. RadGraph-XL addresses this by expanding knowledge graph construction to other medical imaging modalities and anatomies. This expansion facilitates the development of general medical knowledge databases.
@@ 9.2 Evaluation of medical correctness
Evaluating the medical correctness of generated reports is crucial for clinical applications. Compared to previous works (Messina et al., 2022; Liao et al., 2023), recent works have paid more attention to it, but still have two shortcomings. First, in the reviewed articles, medical correctness evaluation has only been applied to chest X-ray reports. Second, the evaluation is based on the automatic labeler of radiology reports, resulting in a limited scope and reliability of the labels. For instance, the labeler in the medical efficacy metric only targeted 14 types of diseases and the average F1 score is around 0.798 (Smit et al., 2020a). Improving the accuracy and scale of automatic labeling tools can help optimize the evaluation process.
9.3 Large public datasets and unified benchmark comparison
@@
As shown in Table 2, most public datasets are limited in size. Deep learning-based techniques require a large amount of data. The contemporary prevalence of large models underscores this need for large data volumes. Among the datasets, the MIMIC dataset and recently released CheXpert Plus are relatively large, but only include chest X-ray data. Large datasets targeting other image modalities and diseases need to be constructed. In addition, while MIMIC-CXR is a well-established benchmark compared to IU-Xray, dataset utilization lacks standardization, complicating comparisons. We urge papers using the MIMIC dataset to define their training, validation, and testing partitions, with explicit disclosure of the filtration method, particularly for the testing dataset. Open-sourcing the code can also facilitate the comparison of different methods on a unified test set. The recently introduced ReXrank leaderboard significantly addresses benchmark inconsistency, marking a milestone contribution despite focusing only on chest X-rays.
9.4 Human-AI interaction and explainability
Effective human-computer interaction can improve the explainability of models. Recent studies have increasingly prioritized human-computer interaction. For traditional deep learning-based methods, Tanwani et al. (2022) constructed a visual question-answering system for the generation of medical reports, and Tanida et al. (2023) linked the output results to image regions using object detection, allowing users to select areas of interest and receive the corresponding language explanations. However, these systems lack comprehensive language understanding, limiting meaningful interaction with users. The advent of large models with advanced language comprehension presents new opportunities. General dialogue systems, such as GPT-4, have demonstrated the ability to articulate their thought processes and provide rigorous logical answers. In contrast, specialized large models for report generation have yet to achieve comparable advancements. Eight recent studies have explored report-generation tasks using large models. One of these studies produces additional bounding boxes to indicate the areas targeted by the generated descriptions (Bannur et al., 2024). However, the design of prompts is still relatively uniform, and none have tested whether these models can articulate their thought processes by responding to queries such as “How did you determine the location of the lesion?” Additionally, while some research has incorporated patient clinical histories at the input stage, there has been no exploration of scenarios where users might wish to provide medical history in queries, for example “I previously had tuberculosis. What impact would that have?"
Providing quantitative results can also improve the explainability of the system and increase user trust. Although measuring uncertainty is prevalent in medical image analysis (Lambert et al., 2024), only one reviewed paper (Wang et al., 2024b) offered quantitative assessments of system uncertainty, suggesting a promising avenue for future research.
@@ 9.5 Standardized report generation
Most existing methods primarily address unstructured report generation. In contrast, structured reporting offers several advantages, including time savings (Hong and Kahn, 2013; Nobel et al., 2022), error reduction, lower communication costs associated with ambiguous language, and improved data consistency and comparability for the development of data records and research. Pellegrini et al. (2023a) developed a structured template and released a related dataset. This promising start could pave the way for further exploration in this direction. It is noteworthy that effective template design is essential for structural reporting; poorly designed templates can fragment related processes into subcategories, disrupt cohesive synthesis, and hinder comprehensive understanding, especially in multi-compartmental pathological conditions. Currently, research on designing effective structured report templates is limited, making it a valuable area for investigation.
@@ 10 Conclusions
In this study, we have conducted a detailed technical review of 100 papers on automatic medical report generation published in the years 2021, 2022, 2023, and 2024 to showcase both mainstream and novel techniques. Our particular focus lies on the utilization and fusion of multi-modality data. The analysis of methods is structured based on the components of the report generation pipeline, presenting the key techniques for each component. Moreover, we examine recent advancements in large model applications and model explainability and summarize current publicly available datasets and evaluation methods, including both quantitative and qualitative assessments. Subsequently, we compare the results of a subset of the reviewed articles under the same experimental setting. Finally, we outline the current challenges and propose future directions for the generation of medical reports. In general, sustained progress is needed to produce high-quality reports. This survey aims to offer a comprehensive overview of report generation techniques, emphasize critical issues, and assist researchers in quickly grasping recent advances in the field to build stronger systems for clinical practice.
@@
Table A1: Summary of papers in the survey. The overview is based on the findings of the survey. The dataset and metrics section focuses on mainstream datasets and metrics. ’–’ indicates that the paper either does not detail this process or does not include the key techniques summarized in this survey. The following abbreviations are used: I-Archi: the architecture of image feature learning, I-Module: the enhancement module of image feature learning, NI: the feature learning of non-image data, GEI: gastrointestinal endoscope image, RetiI: retinal image, Term: terminology, KnowB: knowledge base, RealR: real report, ClinicalI: clinical information, InsT: instructions, FreFilter: frequency-based filtering, ConAtoL: converting all tokens to lowercase, RemoNAT: removing non-alphabetic tokens, DenN: DenseNet, IncepV3: InceptionV3, EffNet: EfficientNet, TranF: Transformer, AT: auxiliary task, ContL: Contrastive learning, ContM: Contrastive module, Embed: general embedding techniques, MM: memory metric, GraA: graph attention, AttenT: attention, FeatO: feature-level operation, OptimS: optimization strategies, H-LSTM: hierarchical LSTM, ReLoss: re-weighted loss function, ReinL: reinforcement learning, R-L: Rouge-L, MET: METEOR, C-D: CIDEr-D, CE: clinical efficacy, R&T: retrieval-based and template-based techniques, RGF1: RadGraphF1, Com: comparison, Clas: classification, Escore: error scoring, ECI: extracting case-related information, IU: IU-Xray, MIMIC: MIMIC-CXR, ImaGeno: Chest ImaGenome, COV: COV-CTR, DeepEye: DeepEyeNet, CCT: COVID-19 CT, Ret-I: Retina ImBank, and Ret-C: Retina Chinese. In addition, AT-Graph, AT-Class, AT-EC, and AT-DS mean the graph-based, classification, embedding comparison, and detection/segmentation auxiliary tasks.
Paper Input Data Data Preparation Feature Learning Feature Fusion Generation Datasets Metrics
Image Non-image Decoder-based method R&T
Archi A_L E_M Archi E_M Archi T_S
Traditional deep learning methods
@@
ResNet AT-Graph Embed,
TranF
FeatO TranF IU,
@@
Liu et al. [2021d] Chest X-ray ResNet AT-Graph TranF IU,
@@
CNNs LSTMs Curriculum learning IU,
@@
You et al. [2021] Chest X-ray ResNet AttenT TranF IU,
@@
Miura et al. [2021] Chest X-ray DenN+
TranF
TranF ReinL MIMIC BLEU,
@@
Alfarghaly et al. [2021] Chest X-ray Resizing DenN AT-Class AttenT TranF IU BLEU,
@@
Embed
AttenT H-LSTM IU,
@@
CNNs Embed FeatO LSTM DeepEye BLEU,
@@
DenN AT-Class Embed,
BERT
FeatO,
AttenT
TranF AT CCT BLEU,
@@
Endo et al. [2021] Chest X-ray ResNet R MIMIC BLEU,
@@
ResNet AT-Class AttenT H-LSTM ReinL IU,
@@
Pino et al. [2021] Chest X-ray DenN T IU,
@@
ResNet ContM H-LSTM IU,
@@
DenN TranF MIMIC BLEU,
@@
CNNs Embed LSTM LSTM DeepEye BLEU,
@@
IncepV3 AT-Class LSTM IU,
@@
FreFilter,
ECI
ResNet Embed,
RotatE,
@@
DenN+
TranF
AT-EC TranF,
@@
DenN AT-Graph H-LSTM IU BLEU,
@@
TranF
AT-Graph TranF FFA-IR BLEU,
@@
TranF
MM TranF ReinL IU,
@@
TranF
ContL MM TranF IU,
@@
DenN+
TranF
AT-Graph Embed,
BERT
FeatO,
AttenT
TranF IU BLEU,
@@
TranF TranF IU BLEU,
@@
Yan [2022] Chest X-ray DenN+
TranF
AT-Graph MM TranF IU,
@@
TranF
AT-Class TranF AT IU BLEU,
@@
TranF
AT-Class AttenT,
@@
MET
Wang et al. [2022d] Chest X-ray Resizing TranF AT-Class,
@@
CNNs Embed AttenT LSTM DeepEye BLEU,
@@
DenN ContM FeatO,
AttenT
TranF IU,
@@
Jia et al. [2022] Chest X-ray DenN AttenT H-LSTM IU,
@@
Zhang et al. [2022] Chest X-ray Resizing VGG R MIMIC Precision
@@
TranF
AT-Graph Embed AttenT TranF MIMIC BLEU,
@@
TranF
AT AttenT TranF IU,
@@
DenN AT-Class LSTM ReLoss IU BLEU,
@@
MET
Abela et al. [2022] Chest X-ray DenN T MIMIC BLEU,
@@
Wu et al. [2022] Chest X-ray ResNet ContL LSTM IU,
@@
ContL
Embed,
BERT
AttenT TranF IU BLEU
@@
MET
Kong et al. [2022] Chest X-ray Resizing TranF R IU,
@@
TranF AT-Class,
ContL
FeatO TranF MIMIC BLEU,
@@
Wang et al. [2023d] Chest X-ray Tokenizing TranF AT TranF IU,
@@
DenN AT-Graph AttenT TranF IU,
@@
TranF
TranF ReinL,
@@
Tokenizing,
ECI
TranF AT-Class,
ContL
Embed,
@@
GraA FeatO,
AttenT
TranF IU,
@@
Zhang et al. [2023b] Chest X-ray DenN+
TranF
@@
ECI DenN+
TranF
AT-Class Embed,
@@
TranF
ContL TranF IU,
@@
DenN+
TranF
Embed,
BERT
FeatO,
AttenT,
@@
DenN ContL Embed,
TranF
FeatO,
AttenT
TranF IU,
@@
Pellegrini et al. [2023a] Chest X-ray,
InsT
EffNet Embed,
BERT
AttenT Classification for structural report generation Rad-ReStruct CE
@@
AT-DS Embed AttenT TranF MIMIC,
@@
Rewei-ghting AttenT TranF - IU,
@@
Embed AttenT TranF - MIMIC,
@@
Hou et al. [2023] Chest X-ray ResNet+
TranF
AT-Class TranF IU,
MIMIC
BLEU,
R-L,
MET,
CE
Jeong et al. [2024] Chest X-ray Resizing TranF R MIMIC BLEU,
Escore,
RGF1
@@
Embed,
TranF
FeatO,
AttenT
TranF ReinL IU,
@@
TranF
AT TranF ReLoss IU,
@@
TranF
Embed,
TranF
@@
ResNet AT-Class Embed,
TranF
FeatO,
AttenT
TranF IU,
@@
Large model-based methods
Pellegrini et al. [2023b] Chest X-ray,
InsT
ResNet+
TranF
Embed FeatO,
AttenT
TranF MIMIC BLEU,
R-L,
MET,
CE
Lee et al. [2023] Chest X-ray,
InsT
CNNs Embed FeatO,
AttenT
TranF MIMIC BLEU,
R-L,
MET,
C-D,
CE
Wang et al. [2023e] Chest X-ray,
InsT
TranF Embed FeatO,
AttenT
TranF IU,
MIMIC
BLEU,
R-L,
MET,
C-D,
CE
Hyland et al. [2023] Chest X-ray,
InsT,
ClinicalI
TranF Embed FeatO,
AttenT
TranF MIMIC BLEU,
R-L,
MET,
CE,
RGF1
Bannur et al. [2024] Chest X-ray,
InsT,
ClinicalI,
RealR
TranF Embed FeatO,
AttenT
TranF IU,
MIMIC
BLEU,
R-L,
MET,
CE,
RGF1
Chen et al. [2024] Chest X-ray,
InsT
TranF Embed FeatO,
AttenT
TranF MIMIC RG-based,
GPT-4-based,
Grading
Guo et al. [2024] Ultrasound,
InsT
TranF Embed FeatO,
AttenT
TranF
Zhang et al. [2024a] Chest X-ray,
InsT
ResNet Embed FeatO,
AttenT
TranF IU,
MIMIC
Human-based evaluation
Li et al. [2024] Chest X-ray,
InsT
TranF Embed FeatO,
AttenT
TranF GPT-4-based
Alkhaldi et al. [2024] Chest X-ray,
InsT
TranF Embed FeatO,
AttenT
TranF MIMIC
Wu et al. [2023a] Chest X-ray,
InsT
TranF Embed FeatO,
AttenT
TranF IU,
MIMIC
BLEU,
R-L,
CE
Zhou et al. [2024] Chest X-ray,
InsT
TranF Embed FeatO,
AttenT
TranF IU,
MIMIC
BLEU, CE, RGF
Table A2: Comparisons of the model performance on the MIMIC-CXR Dataset. B1, B2, B3, B4, R-L, C-D, P, R, and F represent BLEU-1, BLEU-2, BLEU-3, BLEU-4, ROUGE-L, CIDEr-D, precision, recall, and F1 score, respectively. The best and second best results are highlighted. All values were extracted from their papers. * means large model-based methods.